OpenStudy (anonymous):

How much voltage is required to cause 0.003 amps of current to flow through a 5000 omh resistor?

OpenStudy (amistre64):

theres usually an equation that goes with this

OpenStudy (amistre64):

v = a*o

OpenStudy (amistre64):

id guess; 15 volts

OpenStudy (radar):

Use I = e/R \[.003=e/r\]

OpenStudy (radar):

3 mils at 5K is just about 15 volts just as amistre64 guessed.

OpenStudy (anonymous):

will you please show your work

OpenStudy (radar):

e=ir \[e=i \times r = .003Amperes \times 5000\Omega\] e=15 volts.

OpenStudy (radar):

3 10^-3 times 5 10^3= 15 volts.